4 pipes take 70 minutes to fill a tank. how long will it take for 7 pipes to fill the tank?
Answer by TimothyLamb(4379)

r = w/pt
1/(4*70) = 1/(7*t)
1/280 = 1/7t
t = (1/7)/(1/280)
t = 40 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php